# Quantised helicity in optical media

### v1

@@ Abstract
$-i\Lambda\left[\hat{a}_{\lambda}(\mathbf{k})-\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{b}_{\lambda}(-\mathbf{k})+\hat{b}^{\dagger}_{\lambda}(\mathbf{k})\right]$
$-\lambda\Lambda\left[\hat{a}_{\lambda}(\mathbf{k})+\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{c}_{\lambda}(-\mathbf{k})+\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\right].$
where $\hat{a}_{\lambda}(\mathbf{k})$, $\hat{b}_{\lambda}(\mathbf{k})$ and $\hat{c}_{\lambda}(\mathbf{k})$ are the annihilation operators corresponding to the fields $\mathbf{A}$, $\mathbf{P}$ and $\mathbf{M}$. We have also defined $\widetilde{\omega}_{0}^{2}=\omega_{0}^{2}+\alpha^{2}$, as well as the coupling parameter $\Lambda=(\alpha/2)\sqrt{k/\widetilde{\omega}_{0}}$. We define the circular polarisation vectors $\mathbf{e}_{\pm}$ with the convention that that $\mathbf{e}_{\lambda}(\mathbf{k})\cdot\mathbf{e}_{\lambda^{\prime}}(\mathbf{k})=\delta_{\lambda\lambda^{\prime}}$, $\mathbf{e}_{\lambda}(\mathbf{k})\cdot\mathbf{k}=0$ for all $\lambda$, and $\mathbf{e}_{\lambda}(-\mathbf{k})=-\mathbf{e}_{-\lambda}(\mathbf{k})$. Our continuum annihilation and creation operators satisfy the bosonic commutation relations $\left[\hat{a}_{\lambda}(\mathbf{k}),\hat{a}^{\dagger}_{\lambda^{\prime}}(\mathbf{k}^{\prime})\right]=\delta_{\lambda\lambda^{\prime}}\delta(\mathbf{k}-\mathbf{k}^{\prime})$, with similar relations for $\hat{b}_{\lambda}(\mathbf{k})$ and $\hat{c}_{\lambda}(\mathbf{k})$, and all other commutators equal to zero. The explicit expansions of $\hat{\mathbf{A}}$, $\hat{\mathbf{P}}$ and $\hat{\mathbf{M}}$ used to obtain (22) from (18) are given in the Supplemental Material [24].
To diagonalise (22), we follow [22] and define polariton annihilation operators as linear combinations of the $\mathbf{A}$, $\mathbf{P}$ and $\mathbf{M}$ creation/annihilation operators,
@@
with the requirement that $\left[\hat{p}_{i}(\mathbf{k},\lambda),\hat{H}\right]=\omega_{i}\hat{p}_{i}(\mathbf{k},\lambda)$. The coefficients $\alpha_{i}$, $\widetilde{\alpha}_{i}$, $\beta_{i}$, $\widetilde{\beta}_{i}$, $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ are constants, and the index $i$ labels the different polariton branches. The details of the diagonalisation are given in the supplemental material [24]. We find that the polariton branch frequencies, $\omega_{i}$, satisfy the dispersion relation
$k^{2}=\omega_{i}^{2}n^{2}(\omega_{i})=\omega_{i}^{2}\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right)^{2}$ (24)
the solutions of which are plotted in Fig. 2.
Figure 2: The polariton branches, i.e. solutions to the dispersion relation in Eq. (24), with $\alpha=\omega_{0}/2$, chosen to ensure that the wave-vector and group velocity direction coincide.
@@
$:\widehat{\mathcal{H}}:$ $=\int d^{3}r\;:\frac{1}{2}\left(\mathbf{\widehat{A}}\cdot\mathbf{\widehat{B}}-\mathbf{\widehat{C}}\cdot\mathbf{\widehat{D}}\right)+\boldsymbol{\widehat{\Pi}}^{P}\cdot\mathbf{\widehat{M}}-\boldsymbol{\widehat{\Pi}}^{M}\cdot\mathbf{\widehat{P}}:$
$=\sum_{\lambda=\pm}\int d^{3}k\;\lambda\;\bigg{[}\hat{p}^{\dagger}_{u}(\mathbf{k},\lambda)\hat{p}_{u}(\mathbf{k},\lambda)+\hat{p}^{\dagger}_{d}(\mathbf{k},\lambda)\hat{p}_{d}(\mathbf{k},\lambda)$
$\hskip 85.35826pt-\hat{p}^{\dagger}_{p}(\mathbf{k},\lambda)\hat{p}_{p}(\mathbf{k},\lambda)\bigg{]},$ (25)
@@
where we have normal-ordered with respect to the polariton vacuum, and have labelled the three polariton branches $u$, $d$ and $p$. Equation (Quantised helicity in optical media) is simply the difference between the number of positive and negative helicity polaritons, and is completely analogous to the free space expression (21). Note that one of the modes has a negative refractive index, and therefore contributes to the total helicity with opposite sign (as for this mode, the wave propagates in the opposite direction to the wave vector).
@@
We note that the helicity is only expressible in this simple way when the medium is dual-symmetric (i.e. when $\alpha=\beta$ and $\omega_{E}=\omega_{M}$). This is to be expected as the helicity is only expressible in terms of polariton number operators in a dual-symmetric medium, as in a general lossless (but non-dual-symmetric) medium, the number of polaritons is conserved while the helicity is not. In such a medium, the total helicity operator can still be expressed in terms of circularly polarised polaritons, but will include more general products of polariton creation and annihilation operators.

### v2

@@ V.1 Helicity in dual symmetric media
$-i\kappa\left[\hat{a}_{\lambda}(\mathbf{k})-\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{b}_{\lambda}(-\mathbf{k})+\hat{b}^{\dagger}_{\lambda}(\mathbf{k})\right]$
$-\lambda\kappa\left[\hat{a}_{\lambda}(\mathbf{k})+\hat{a}^{\dagger}_{\lambda}(-\mathbf{k})\right]\left[\hat{c}_{\lambda}(-\mathbf{k})+\hat{c}^{\dagger}_{\lambda}(\mathbf{k})\right].$
where we have defined the coupling parameter $\kappa=(\alpha/2)\sqrt{k/\widetilde{\omega}_{0}}$. To diagonalise (26), we follow [25, 30] and define polariton annihilation operators as linear combinations of the $\mathbf{A}$, $\mathbf{P}$ and $\mathbf{M}$ creation/annihilation operators
@@
with the requirement that $\hat{p}_{i}(\mathbf{k},\lambda),\hat{H}]=\omega_{i}\hat{p}_{i}(\mathbf{k},\lambda)$. The coefficients $\alpha_{i}$, $\widetilde{\alpha}_{i}$, $\beta_{i}$, $\widetilde{\beta}_{i}$, $\gamma_{i}$ and $\widetilde{\gamma}_{i}$ are constants, and the index $i$ labels the different polariton branches. The details of the diagonalisation are given in Appendix C. We find that the polariton branch frequencies, $\omega_{i}$, satisfy the dispersion relation
$k^{2}=\omega_{i}^{2}n^{2}(\omega_{i})=\omega_{i}^{2}\left(1-\frac{\alpha^{2}}{\omega_{i}^{2}-\omega_{0}^{2}}\right)^{2}.$ (28)
The solutions to Eq. (28) are plotted in Fig. 2.
Figure 2: The polariton branches, i.e. solutions to the dispersion relation in Eq. (28), with $\alpha=\omega_{0}/2$. The sign of $\omega_{i}$ has been chosen to ensure that the wave-vector and group velocity direction coincide.
Figure 3: Some illustrative examples of the expectation value of $\hat{h}$ for an excitation in a single polariton mode, with a range of non-dual-symmetric model parameters. The corresponding dispersion curves for the polariton modes are shown below each graph. Figure 3a shows a typical case, with the helicity approaching unity when the branches are photon-like. Note that $\langle\hat{h}\rangle$ is not bounded between $\pm 1$, and the middle branch’s contribution changes sign as $k$ varies. Figure 3b shows a case with $\omega_{E}=\omega_{M}$, but $\alpha\neq\beta$, where the asymptotic behaviour of the matter-like branches does not approach 0 or 11 for large $k$. Figure 3c shows a nonmagnetic material, which can be described in this formalism by letting $\omega_{M}\rightarrow\infty$. One of the branches then becomes a flat band at a very high frequency (not shown), and the system effectively reduces to two branches.
@@
[…]
$\hskip 85.35826pt+\boldsymbol{\widehat{\Pi}}^{P}\cdot\mathbf{\widehat{M}}-\boldsymbol{\widehat{\Pi}}^{M}\cdot\mathbf{\widehat{P}}$:
$=\sum_{\lambda=\pm}\int d^{3}k\;\lambda\;\bigg[\hat{p}^{\dagger}_{u}(\mathbf{k},\lambda)\hat{p}_{u}(\mathbf{k},\lambda)$ (30)
$\hskip 85.35826pt+\hat{p}^{\dagger}_{d}(\mathbf{k},\lambda)\hat{p}_{d}(\mathbf{k},\lambda)$
$\hskip 85.35826pt-\hat{p}^{\dagger}_{p}(\mathbf{k},\lambda)\hat{p}_{p}(\mathbf{k},\lambda)\bigg],$
@@
where we have normal-ordered with respect to the polariton vacuum, and have labelled the three polariton branches $u$, $d$ and $p$. More details of this result’s derivation are presented in Appendix D. Equation (29) is simply the difference between the number of positive and negative helicity polaritons, and is completely analogous to the free space expression (24). As polaritons are the eigenstates of the full Hamiltonian, the helicity equals $\pm$ the total energy density divided by frequency, rather than just the electromagnetic energy density. Note that one of the modes has a negative refractive index, and therefore contributes to the total helicity with opposite sign (as for this mode, the wave propagates in the opposite direction to the wave vector).
@@
We note that the helicity is only expressible in this simple way when the medium is dual-symmetric (i.e. when $\alpha=\beta$ and $\omega_{E}=\omega_{M}$ such that $\varepsilon/\mu=1$). This is to be expected, as the helicity is only expressible in terms of polariton number operators in a dual-symmetric medium. In a general lossless medium, the number of polaritons is conserved while the helicity is not. In such a medium, the total helicity operator can still be expressed in terms of circularly polarised polaritons, but includes more general products of polariton creation and annihilation operators, given in Appendix D (D). In the following subsection, we will discuss the helicity in such cases. The helicity and the Hamiltonian no longer commute, and we show that the expectation value of the helicity acquires an explicit time-dependence.
V.2 Helicity in ordinary media
The dual symmetric Hamiltonian considered in the previous section represents a very special case, where the helicity is exactly conserved. When $\alpha\neq\beta$ and/or $\omega_{E}\neq\omega_{M}$, the transformation (II) is no longer a symmetry of the Hamiltonian (25), and the helicity is no longer locally conserved. A consequence of this is that the commutator $\hat{H},\int d^{3}r\;{\hat{h}}]\neq 0$, meaning that eigenstates of the Hamiltonian are not helicity eigenstates. Explicitly, the commutator is given by
$\left[\hat{H},\int d^{3}r\;{\hat{h}}\right$] $=\sum_{\lambda}\frac{i}{2}\Bigg(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}+\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Bigg)(\tilde{\omega}_{E}-\tilde{\omega}_{M})\hat{b}_{\lambda}^{\dagger}(\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})+\frac{i}{2}\Bigg(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}-\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Bigg)(\tilde{\omega}_{E}+\tilde{\omega}_{M})\hat{b}_{\lambda}(-\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})$
$\hskip 14.22636pt+\frac{i\lambda}{2}\sqrt{\frac{k}{\tilde{\omega}_{E}}}(\alpha-\beta)\Big(-\hat{a}_{\lambda}(-\mathbf{k})\hat{b}_{\lambda}(\mathbf{k})-\hat{b}_{\lambda}(\mathbf{k})\hat{a}_{\lambda}^{\dagger}(\mathbf{k})\Big)+\frac{1}{2}\sqrt{\frac{k}{\tilde{\omega}_{M}}}(\alpha-\beta)\Big(\hat{a}_{\lambda}(-\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})-\hat{a}_{\lambda}^{\dagger}(\mathbf{k})\hat{c}_{\lambda}(\mathbf{k})\Big)$
$\hskip 14.22636pt-\text{H.c.},$ (31)
which is generally non-zero. To arrive at this result, we have expressed $\hat{h}$ in terms of the $\hat{a}$, $\hat{b}$ and $\hat{c}$ operators, which can be found in Appendix D (D).
While the commutator (31) is generally non-zero, it can be shown that its expectation value is equal to zero in the case of a polariton number state. For lossless media, polariton number states therefore exhibit a time-independent mean helicity. However, the value of this mean helicity is not generally $\pm 1$ for each circularly polarised polariton. This follows directly from the fact that we included the contribution associated with the matter in the helicity, which is not necessarily closely tied to the polariton circular polarisations. For a state with $n$ polaritons of wavevector $\mathbf{k}$ in branch $i$, we have
$\int d^{3}r\;$ $\left\langle{n(i,\mathbf{k},\lambda)}\right|\hat{h}(\mathbf{r})\left|{n(i,\mathbf{k},\lambda)}\right\rangle=$
$n\lambda(\alpha_{i}\alpha_{i}^{*}+\tilde{\alpha}_{i}\tilde{\alpha}_{i}^{*})+\Bigg(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}+\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Bigg)\mathrm{Im}[\tilde{\beta}_{i}\tilde{\gamma}_{i}^{*}-\beta_{i}\gamma_{i}^{*}]+\Bigg(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}-\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Bigg)\mathrm{Im}[\beta_{i}\tilde{\gamma}_{i}^{*}-\tilde{\beta}_{i}\gamma_{i}^{*}],$ (32)
with no sum over the repeated index $i$ implied. Here, $\beta_{i}$, $\gamma_{i}$, $\tilde{\beta}_{i}$ and $\tilde{\gamma}_{i}$ now refer to the polariton expansion coefficients that diagonalise the general Hamiltonian (25), rather than the dual-symmetric Hamiltonian (26). Explicit expressions for these are given in Appendix E.
Plots of the expectation value (32) are shown in Figure 3 for various combinations of model parameters, along with the corresponding polariton dispersion curves. Unlike in the dual-symmetric case, the total helicity of a polariton number state $\left|{n(i,\mathbf{k},\lambda)}\right\rangle$ is now $k$ dependent. It reaches a constant value only at large $k$, where we recover the free-space behaviour of one photon-like branch with unit helicity. This behaviour is to be expected, as the upper polariton branch will become increasingly photon-like at high $k$, where the helicity must be $\pm 1$ for the circular polarisations. In Figure 3a we show the helicity for an example medium where $\omega_{E}\neq\omega_{M}$ and $\alpha\neq\beta$, and we see that the branches typically have significant helicity only when they are photon-like.
When the resonant frequencies $\omega_{E}$ and $\omega_{M}$ are equal, illustrated in Figure 3b, we find that the photon-like branch still behaves in this manner. However, we also see that the matter-like branches carry significant helicity. Matching the resonant frequencies, but not the coupling strengths, therefore retains some features of the fully dual-symmetric case: the three branches asymptotically have constant, non-zero helicity. However, unlike the dual-symmetric case, the helicity reaches this value only asymptotically as $k$ increases. Furthermore, the asymptotic helicity values are not $\pm 1$ when $\alpha\neq\beta$. On a mathematical level, this is due to the factors of $\alpha/\beta$ and $\beta/\alpha$ in Eq. (32). We therefore conclude that when the polarisation and magnetisation respond in phase, but at different strengths, enough symmetry is retained for the matter-like polaritons to carry helicity.
A special case of practical interest is the description of a non-magnetic medium. Although the formalism for defining helicity presented here requires that the material possesses both an electric and a magnetic response, a non-magnetic limit can be taken by letting $\omega_{M}\rightarrow\infty$, as shown in Figure 3c. This effectively reduces the system to two branches, as one branch becomes a flat band at very high frequency. Nevertheless, the operator (9), obtained from the duality rotation (II), may still be used to define the helicity. Note that for a polariton in the high frequency branch, $\langle\hat{h}\rangle\sim 0$.
V.3 Helicity oscillations
While the polariton number states considered above have a time-independent mean helicity, the helicity does generally depend on time. In this subsection, we explore the time-dependence of the helicity, and demonstrate that the average helicity is not always conserved through an explicit example. For this, we consider preparing a system in an eigenstate of helicity, and we calculate the expectation value of the total helicity as the state evolves in time. In order to construct a helicity eigenstate, we note that the helicity operator can be diagonalised by a similar procedure to that which we used to diagonalise the Hamiltonian in section V.1. We write a linear combination of polariton creation annihilation operators,
$\hat{\mathfrak{h}}(\mathbf{k},\lambda)=\sum_{i}a_{i}\hat{p}_{i}(\mathbf{k},\lambda)+\tilde{a}_{i}\hat{p}^{\dagger}_{i}(-\mathbf{k},\lambda)$ (33)
and impose the condition that $\int d^{3}r\,\hat{h},\hat{\mathfrak{h}}]=\Lambda\hat{\mathfrak{h}}$. The operator $\hat{\mathfrak{h}}$ may then be interpreted as an annihilation operator for a mode of well-defined helicity with helicity eigenvalue $\Lambda$. We note that the helicity eigenstates do not generally have eigenvalues of $\pm 1$, due to the addition of the matter component.
Helicity eigenstates can be explicitly constructed by applying the creation operator $\hat{\mathfrak{h}}^{\dagger}(\mathbf{k},\lambda)$ to the vacuum state. However, the helicity vacuum is generally squeezed with respect to the polariton vacuum (as the helicty vacuum is defined through a mixture of polariton annihilation and creation operators). This is difficult to realise experimentally, and we will therefore proceed by approximating the helicity eigenstate using only the polariton-number-conserving terms in the total helicity $\int d^{3}r\,\hat{h}$. This is equivalent to applying $\hat{\mathfrak{h}}^{\dagger}(\mathbf{k},\lambda)$ to the polariton vacuum. In other words, we excite a mixture of all polariton branches at a single wavenumber. We note that this approximation does not impact the qualitative points that we wish to illustrate, and is a good approximation given that we are considering wavenumbers away from the medium resonances.
In this approximation, the helicity creation operator acting on the vacuum can be written $\hat{\mathfrak{h}}^{\dagger}(\mathbf{k},\lambda)\left|{0_{\text{polariton}}}\right\rangle=\sum_{i}a_{i}\hat{p}^{\dagger}(\mathbf{k},\lambda)\left|{0_{\text{polariton}}}\right\rangle$, as $\hat{p}_{i}\left|{0_{\text{polariton}}}\right\rangle=0$. This means that the terms proportional to $\hat{p}_{i}$ in $\hat{\mathfrak{h}}^{\dagger}(\mathbf{k},\lambda)$ (or equivalently, those proportional to $\hat{p}^{\dagger}_{i}$ in Eq. (33)) will not contribute to the construction of a helicity eigenstate. We therefore see that we only need to construct an eigenstate of the number-conserving part of the helicity operator. The full helicity operator is given in terms of the polariton operators in Appendix D (D), and the number-conserving part of this is given by
$\int d^{3}r\,\hat{h}^{\text{num}}(t)=\int d^{3}k\;\sum_{ij,\,\lambda}h_{ij}(\lambda)\hat{p}^{\dagger}_{i}(\mathbf{k},\lambda)\hat{p}_{j}(\mathbf{k},\lambda),$ (34)
where we have introduced the shorthand
$h_{ij}(\lambda)=\lambda\alpha_{i}\alpha^{*}_{j}+\lambda\tilde{\alpha}^{*}_{j}\tilde{\alpha}_{i}$ (35)
$+\frac{i}{2}\Big(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}+\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Big)(\gamma^{*}_{j}\beta_{i}+\tilde{\gamma}_{i}\tilde{\beta}_{j}^{*}-\gamma_{i}\beta_{j}^{*}-\tilde{\gamma}_{j}^{*}\tilde{\beta}_{i})$
$-\frac{i}{2}\Big(\frac{\beta}{\alpha}\sqrt{\frac{\tilde{\omega}_{E}}{\tilde{\omega}_{M}}}-\frac{\alpha}{\beta}\sqrt{\frac{\tilde{\omega}_{M}}{\tilde{\omega}_{E}}}\Big)(-\tilde{\beta}_{i}\gamma_{j}^{*}-\beta_{j}^{*}\tilde{\gamma}_{i}+\beta_{i}\tilde{\gamma}_{j}^{*}+\tilde{\beta}_{j}^{*}\gamma_{i}).$
The eigenvectors of the matrix $h_{ij}$ contain the required coefficients to write down a state $\left|{\psi}\right\rangle$, such that $\int d^{3}r\,\hat{h}\left|{\psi}\right\rangle=\Lambda\left|{\psi}\right\rangle$, in terms of a superposition of the polariton operators. This generally takes the form
$\left|{\psi}\right\rangle=\left[a_{1}\hat{p}^{\dagger}_{1}(\mathbf{k},\lambda)+a_{2}\hat{p}^{\dagger}_{2}(\mathbf{k},\lambda)+a_{3}\hat{p}^{\dagger}_{3}(\mathbf{k},\lambda)\right]\left|{0}\right\rangle,$ (36)
where $a_{i}$ are the coefficients for the different polariton branches, normalised such that $|a_{1}|^{2}+|a_{2}|^{2}+|a_{3}|^{2}=1$. As we shall see, the average total helicity will oscillate in time. For this, we should note that, by construction, the time-dependence of the polariton operators is $\hat{p}_{i}(\mathbf{k},\lambda,t)=\hat{p}_{i}(\mathbf{k},\lambda)\exp(-i\omega_{i}t)$. The total helicity density therefore takes the form
$\int d^{3}r\,\hat{h}(t)$ $=\int d^{3}k\;\sum_{ij,\,\lambda}h_{ij}(\lambda)\hat{p}^{\dagger}_{i}(\mathbf{k},\lambda)\hat{p}_{j}(\mathbf{k},\lambda)e^{i\left(\omega_{i}-\omega_{j}\right)t}$
$+\text{(non-polariton-number-conserving terms)},$ (37)
where we have not given the explicit form of the non-number conserving terms (i.e. the terms which contain products of two creation or two annihilation operators) as they will vanish when the expectation value of $\hat{h}$ is taken. We can explicitly write down the time-dependence of the helicity by taking the expectation value of the operator (34) in the state (36). Defining phases $\phi_{i}$ for the coefficients in (36), so that $a_{i}=|a_{i}|\exp(i\phi_{i})$, we see that
$\int d^{3}r\,\left\langle{\psi}\right|\hat{h}(t)\left|{\psi}\right\rangle$ $=\sum_{\lambda^{\prime}}\int d^{3}k^{\prime}\;\sum_{ij}h_{ij}(\lambda^{\prime})e^{i\left(\omega_{i}-\omega_{j}\right)t}\delta^{3}(\mathbf{k}^{\prime}-\mathbf{k})\delta_{\lambda\lambda^{\prime}}\left[a_{1}^{*}\delta_{i1}+a_{2}^{*}\delta_{i2}+a_{3}^{*}\delta_{i3}\right]\left[a_{1}\delta_{j1}+a_{2}\delta_{j2}+a_{3}\delta_{j3}\right]$
$=\bigg\{|a_{1}|^{2}h_{11}+|a_{2}|^{2}h_{22}+|a_{3}|^{2}h_{33}+2|h_{12}||a_{1}||a_{2}|\cos\left[\left(\omega_{2}-\omega_{1}\right)t+\Delta\phi_{12}\right]$ (38)
$\hskip 56.9055pt+2|h_{13}||a_{1}||a_{3}|\cos\left[\left(\omega_{3}-\omega_{1}\right)t+\Delta\phi_{13}\right]+2|h_{23}||a_{2}||a_{3}|\cos\left[\left(\omega_{3}-\omega_{2}\right)t+\Delta\phi_{23}\right]\bigg\},$
where we have used that $h_{ij}=h_{ji}^{*}$ in the final step, which follows from the fact that $\hat{h}(t)$ is Hermitian, and we have defined $\Delta\phi_{ij}\equiv\phi_{i}-\phi_{j}$. It is evident that the average helicity oscillates in time at the difference frequency between the branches. Only in a dual-symmetric medium are $h_{12}$, $h_{13}$, and $h_{23}$ identically zero, leading to a time-independent average helicity.
An analogy can be drawn between the oscillatory behaviour of the helicity eigenstates shown above, and the phenomenon of neutrino oscillations [31]. A beam of neutrinos prepared in an initially pure flavour eigenstate will be found, upon propagation, to contain a mixture of flavours, as the flavour eigenstates are not eigenstates of the full Hamiltonian [32]. In our case, a similar effect occurs for a system initially prepared in a pure helicity eigenstate, likewise stemming from the fact that the total helicity and the Hamiltonian do not commute. This phenomenon occurs whenever there is a superposition of different polariton branches at the same wavevector $\mathbf{k}$. We note that, in practice, creating this kind of superposition state requires the excitation of both matter-like and photon-like polariton modes.
VI Discussion and conclusion